Compton Meeting, March 15, 2012 (F226)

From PolWiki
Jump to: navigation, search

General

Meeting room: F226 (Game of musical rooms continues..)

Call-in information:

1. Dial Toll-Free Number: 866-740-1260 (U.S. & Canada)
2. International participants dial:
Toll Number: 303-248-0285
Or International Toll-Free Number: http://www.readytalk.com/intl
3. Enter your 7-digit access code, followed by “#”

 --> our access code is 3227471

Agenda

Laser

  1. Moved flipper to position after last lens. Thermal lensing still there.

Photon Detector

  1. LED system: ND filters installed on Delta LED - 20 mV signal but still noisy hclog
  2. LED run 24435 (LED run w/delta at 20 mV). ELOG 183

Electron Detector

  1. More zero crossing analysis: elog 344
  2. Results from 2nd PWTL readout
  3. Updated results plots pre- and post-spot move
  4. Test with detector closer to beam

Miscellaneous

  1. Volunteer for Compton talk at next Wednesday meeting?

Minutes

  1. Flipper will be moved further downstream (again) to see if that helps thermal lensing. Last move didn't seem to help.
  2. Looked at Juan Carlos' elog on LED run 24435. Several questions: if horizontal axis is integrated variable signal, what is size of horizontal error bar? The large variation on Both-Variable (should = delta) does not seem consistent with the plots just prior: analysis bug? Would also be interesting to see distributions.
  3. Looked at Don's photon log on non-zero laser table scintillator differences - totally correlated with photon detector. Wouter found that the FlipPC script had not been run appropriately for this run - so likely had larger than usual position differences, as well as large charge asymmetry.
  4. Vladas discussed his elog entry about finding the "zero-crossing" using a linear fit. The last plot in his entry shows that the difference he sees between the full fit to the spectrum and the linear fit can be attributed to short-comings in the linear fit technique. The red curve shows the results of the same test using pseudo-data.
  5. Dave G. to check with Dipangkar about status of online analysis of parallel PWTL chain.
  6. Vladas will work on updating edet results pre and post-spot move.
  7. Still need to move the edet closer to the beam.
  8. Some discussion of near-term priorities:
    1. Keep Compton running
    2. Understand edet dead time (mostly analysis at this point)
    3. Get LED system running - use regularly.
    4. Wouter comments that there are more systematic tests that should be done. No dis-agreement.